OpenStudy (anonymous):

In a triangle abc,tan a+b/2 cot a-b/2 is equal to

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

is it \[\tan(\frac{a + b}{2}) + \cos(\frac{a - b}{2})\]

OpenStudy (anonymous):

nope cot a-b/2

ganeshie8 (ganeshie8):

\[\tan(\frac{A + B}{2}) + \cot(\frac{A - B}{2})\] like this ?

OpenStudy (anonymous):

tan (a+b)/2*cot(a-b)/2

OpenStudy (anonymous):

\[\tan(A+B)/2*\cot(A+B)/2\]

OpenStudy (anonymous):

sorry cot (a-b)/2

ganeshie8 (ganeshie8):

\[\tan(\frac{A + B}{2}) * \cot(\frac{A - B}{2})\]

ganeshie8 (ganeshie8):

\(\large A + B + C = \pi\)

OpenStudy (anonymous):

c=pi-a+b

OpenStudy (anonymous):

a+b=pi-c

ganeshie8 (ganeshie8):

\[\tan(\frac{A + B}{2}) * \cot(\frac{A - B}{2})\] \[\tan(\frac{C-\pi}{2}) * \cot(\frac{A - B}{2})\] \[\cot(\frac{C}{2}) * \cot(\frac{A - B}{2})\]

OpenStudy (anonymous):

(a - b)/(a + b) = tan [(A-B)/2] / tan [(A+B)/2] (Law of Tangents)

ganeshie8 (ganeshie8):

Oh it equals (a-b)/(a+b) is it ? xD

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so a+b/a-b=cot(a-b)/2*tan(a+b)/2 correct